Line L passes through the points (-2,0) and (0,a). Line LL passes through the points (4,0) and (6,2). What value of a makes the two lines parallel?
A) 1/2
B) -2
C) 2
D) -1/2
E) 0

Answer by mananth(16946)

(-2,0) and (0,a). Line LL passes through the points (4,0) and (6,2).

(-2,0) and (0,a).
slope m = a-0 / 0-(-2)
m=a/2
Line LL passes through the points (4,0) and (6,2).
slope m = 2-0 / 6-4
m=2/2
m= 1
..
for lines to be parallel the slope have to be equal.
a/2 = 1
a= 2
...
m.ananth@hotmail.ca